Solve for x:X e x a_aet=0
Answer #1

Quesi ex.x²-2e²=0 7 ee (x² (z? 2)=0 xx²2 o 7 é But , et to for any XER >x²2=0 =>(x-2)(x+2) = 0 2- TE=0 or at 2=0 Y n = va a=2